OpenStudy (anonymous):

Find all values of c for which the system cx+z=0 2y-4z=0 2x-y=0 has a solution other than (0,0,0) and find all solutions for this value c.

OpenStudy (amistre64):

find the null space

OpenStudy (amistre64):

row reduce echelon the matrix c 0 1 0 2 -4 2 -1 0

OpenStudy (anonymous):

without a value at c how to I find the nul A?

OpenStudy (amistre64):

the value of c tends to migrate along the process

OpenStudy (anonymous):

So c stays c?

OpenStudy (amistre64):

yep, just treat it as some unknown value, other than zero

OpenStudy (amistre64):

or, you can swap rows if need be

OpenStudy (anonymous):

interchange rows to bottom and r1-->r3,

OpenStudy (amistre64):

sure, that sounds like a good step :)

OpenStudy (anonymous):

create vector equation x2=2x3 x1=(1/2)x2

OpenStudy (anonymous):

c=-x3

OpenStudy (amistre64):

2 -1 0 0 2 -4 c 0 1 1 -1/2 0 0 1 -2 c 0 1 1 -1/2 0 0 1 -2 0 -1/2-c -c 1 0 -1 0 1 -2 0 0 -1-2c let -1-2c = 0, c = -1/2 should be one solution

OpenStudy (amistre64):

im sure somethings wrong in that ....

OpenStudy (anonymous):

I had the first portion, but what did you do to get -1/2-c on the reduction

OpenStudy (amistre64):

there it is :) i did addition instead of multiplication

OpenStudy (anonymous):

nevermind I am an idiot

OpenStudy (amistre64):

also, finding a det(A) = 0 seems like a good route

OpenStudy (anonymous):

when it is in reduced form right?

OpenStudy (amistre64):

nah, just take the original form and run a determinant on it, then solve for c -(4+4c+0) c 0 1 c 0 0 2 -4 0 2 2 -1 0 2 -1 +(0+0+0) -4-4c = 0 when c = -1

OpenStudy (amistre64):

\[\begin{pmatrix}x_1&x_2&x_3\end{pmatrix}=x_3\begin{pmatrix}1&2&1\end{pmatrix}\]

OpenStudy (amistre64):

.... latex fail

OpenStudy (amistre64):

the preview isnt previwing :/ \[\begin{pmatrix}x_1\\x_2\\x_3\end{pmatrix}=x_3\begin{pmatrix}1\\2\\1\end{pmatrix}\]

OpenStudy (anonymous):

det A= (c*2*0)+(0*-4*2)+(1*0*1)-((0)+(c*-4*-1)+(1*2*2))=0-4c-4

OpenStudy (amistre64):

yeah, you can go that long route if you like

OpenStudy (anonymous):

c=-1 and then use that in original matrix to find values?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

Ok thanks a ton. This is the easy stuff and I cannot even do it. The worst part is he hard stuff I understand better.

OpenStudy (amistre64):

another route we can tak c 0 1 0 2 -4 2 -1 0 1 0 1/c 0 1 -2 0 -1 -2/c 1 0 1/c 0 1 -2 0 0 (-2c-2)/c c =-1

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Cx+z=0 2y-4z=0 2x-y=0 OR AX=0 Where A=[■(c&0&1@0&2&-4@2&-1&0)],X=[■(x@y@z)] Given system has solutions if |A|=0 |A|=|■(C&0&1@0&2&-4@2&-1&0)|=C|■(2&-4@-1&0)|-0|■(0&-4@2&0)|+1|■(0&2@2&-1)| |A|=C[(2)(0)-(-4)(-1)]-0+1[(0)(-1)-(2)(2)] |A|=C(-4)-4=0 -4C=4 C=4/(-4)=-1 Whenc=-1,given system of equation becomes -x+z=0 2y-4z=0 2x-y=0 Letz=k(≠0) be any number then -x+k=0 orx=k 2y-4k=0 or 2y=4k or y=2k Solution is x=k,y=2k, z=k